OpenStudy (anonymous):

Solve for x. 6 over x equals x over 12

OpenStudy (anonymous):

x = -6 sqrt(2)

OpenStudy (precal):

6/x=x/12 \[x ^{2}=72\] \[x=\pm \sqrt{72}\] \[x=\pm6\sqrt{2}\]